Sharpies Pencils, Inc. has a 4%, semiannual coupon bond with a current market price of $935.50. The bond has a par value of $1,000 and a yield to maturity of 4.6%. How many years is it until this bond matures?
Answer #1

I/Y = 2.3

PMT = 1,000*0.04/2 = 20

PV = -935.5

FV = 1,000

CPT N = 30 payments

No. of years = 30 / 2 payements per year = 15 Years until this bond matures

Can you please upvote? Thank you :-)